OpenStudy (anonymous):

Stewart 7th edition 6.5 number 10

OpenStudy (anonymous):

sorry section 6.5

OpenStudy (dumbcow):

haha please write out the problem :) stewart..hmm sounds like either a calculus textbook or probability

OpenStudy (anonymous):

(a) Find the average value of f on the given interval (b) Find c such that Fave=F(c) (c) Sketch the graph of f and a rectangle whose area is the same as the area under the graph of f. f(x)=1/x, [1,3]

OpenStudy (anonymous):

and yeah it is a calculus textbook haha

OpenStudy (dumbcow):

for avg value within an interval [a,b] \[=\frac{1}{b-a}\int\limits_{a}^{b} f(x) dx\]

OpenStudy (anonymous):

It looks like par a is going to result in 1/2ln(3). Which is 0.54930

OpenStudy (anonymous):

Ok, I got A to be the same, but then how do you start B?

OpenStudy (dumbcow):

well you know the avg f value is ln(3)/2 so: \[f(c) = \frac{1}{c} = \frac{\ln 3}{2}\] solve for c

OpenStudy (anonymous):

Ok, I get .549 to be the answer for c

OpenStudy (anonymous):

Where did f(c) come from?

OpenStudy (dumbcow):

@Enigmatron , its the notation used in the question "c" refers to the x value corresponding to avg function value

OpenStudy (dumbcow):

@mikrol , .549 = ln3/2 which is avg "y" value plus the interval for domain is [1,3]

OpenStudy (anonymous):

@dumbcow I see now. Thank you for the clarification.

OpenStudy (anonymous):

so it would be 2ln/3=c

OpenStudy (dumbcow):

yes.. essentially you take the reciprocal 2/ln3 = 1.82

OpenStudy (anonymous):

Thank you very much that helped a lot!

OpenStudy (dumbcow):

for the rectangle part....length is always size of interval (b-a) and width is always avg function value \[Area = 2*\frac{\ln 3}{2}\]

OpenStudy (anonymous):

So the answer would be 1.82 because we took the reciprocal and not 0.549, correct?

OpenStudy (dumbcow):

yes...since f(x) = 1/x f(1.82) = 1/1.82 = 0.549